A farmer making grape juice fills a glass bottle to the brim and caps it tightly. The juice expands more than the glass when it warms up, in such a way that the volume increases by 0.24%. Calculate the force exerted by the juice per square centimeter if its bulk modulus is 1.8 × 109 N/m2, assuming the bottle does not break.


1
Expert's answer
2021-03-26T11:41:15-0400
"\Delta p=B\frac{\Delta V}{V}\\
\Delta p=1.8\cdot10^9(0.0024)=4.32\cdot10^6\ Pa\\
\Delta p=432\frac{N}{cm^2}"